OpenStudy (anonymous):

A boy have 5 coloured pegs; blue, red, yellow,green and purple. He arranged it randomly in a single horizontal row of 5 pegs. Find the probability the pegs are arranged such that a) blue and yellow occupies both ends b) green is immediately on the left of purple c) blue,red and yellow are together

OpenStudy (anonymous):

(a) P(blue and yellow occupies both ends) = 1/5 x 3/5 x 2/5 x 1/5 x 1/5 = 6/3125 (b) p(green is immediately on the left of purple) = [(1/5x1/5 )x 3/5 x2/5 x1/5 ]x4 =24/3125 (c) P(blue,red and yellow are together) = (3/5 x2/5 x 1/5) x 2/5 x 1/5 x3 = 36/3125 Not sure. Sorry